OpenStudy (anonymous):

Pls help!:) The line x+y=10 intersects the line y=x, the x-axis and the y-axis at the points A, B, C respectively. Find the coordinates of A, B & C

OpenStudy (anonymous):

(10,0),(0,10).(5,5)

OpenStudy (anonymous):

for a, b, and c

OpenStudy (anonymous):

Coordinates of A: \[2x = 10 \rightarrow x=5\] Sub x=5 into x+y=10 (or y=x) to get y=5, so A is (5,5). Coordinates of B: \[x+0=10 \rightarrow x=10\] Sub x=10 into x+y=10 giving y=0, hence B is (10,0) Coordinates of C: \[0+y=10 \rightarrow y=10\] sub y=10 into x+y=10 giving x=0, hence C is (0,10)